[Brahms-l] Shift report 20040314 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sun Mar 14 2004 - 08:11:03 EST
------------------------------------------------------------
Date: Sun Mar 14 08:11:02 2004, Shift: 00:00-08:00
Supervisor: CH, with: 
Log book#15 pages 87 - 87

Shift Summary:
------------------------------------------------------------
1: OK, though we had a very short spill that didn't do 
   anything.

2: 10996:  TPC PEd
   10997:  Pedestal
   10998:  68k tr 2, 88k tr 3
   11001:  BAD RUN
   11002:  38k tr 2, 50k tr 3
   11003:  25k tr 2, 33k tr 3
   11004:  15k tr 2, 20k tr 3

3: D1 tripped (while not taking data - phew!).  
   Called MCR, who sent a guy that did some 
   black magic, mumbled something about opening
   a window, and it worked.  TRMRS seem to have
   lost it's TDCs.  JH is looking at it right now. 
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Mar 14 08:11:23 2004

This archive was generated by hypermail 2.1.8 : Sun Mar 14 2004 - 08:11:45 EST